A bus travelling at a constant velocity of 72km/he passes a stationary car. After a time lapse of 1/6 minutes, the car accelerates uniformly at 3m/s^2 to reach a speed of 108km/hr and maintains this speed afterwards. Draw the speed-time graph and find when and where the car overtakes the bus.


1
Expert's answer
2021-05-11T09:07:29-0400

Velocity of bus, "v_b=72\space km/h=20\space m/s"

Velocity of car, "v_c=108\space km/h=30\space m/s"

Acceleration of car, "a_c=3\space m/s^2"

Acceleration of bus, "a_b=0"

Initial separation of bus and car, "s=20\times10=200\space m"

Velocity of bus with respect to car, "v_{bc}=v_b-v_c=-10\space m/s"

Acceleration of bus with respect to car, "a_{bc}=-3\space m/s^2"

time taken by car to reach bus "=t"

"s=ut+\dfrac{1}{2}at^2"

"-200=-10t+\dfrac{1}{2}(-3)t^2"

"3t^2+20t-200=0"

"t=5.48\space s"